OpenStudy (isabelle99):

Please someone come check my answers Write your salary for the next n years as a geometric sequence in explicit form. What will your salary be in year 5? How much will you have earned in total salary by the end of your 10th year?

OpenStudy (isabelle99):

1st part: \[a _{n}=5000(1+0.05)^{n-1}\] second part: \[a_{5}=5000(1.05)^{5-1}= 6077.53\] And i don't know how to do the last one someone please help

OpenStudy (welshfella):

Use the formula for the sum of n terms of a G S Sn = a1 * (1 - r^n) ------ 1 - r

OpenStudy (welshfella):

a1 = 5000, r = 1.05 and n = 10

OpenStudy (isabelle99):

i got 6300 when i did that. is that right